# Fair Play for Individuals, Foul Play for Groups? Auditing Anonymization’s Impact on ML Fairness

### v1

@@ Abstract
Machine learning (ML) algorithms are heavily based on the availability of training data, which, depending on the domain, often includes sensitive information about data providers. This raises critical privacy concerns. Anonymization techniques have emerged as a practical solution to address these issues by generalizing features or suppressing data to make it more difficult to accurately identify individuals. Although recent studies have shown that privacy-enhancing technologies can influence ML predictions across different subgroups, thus affecting fair decision-making, the specific effects of anonymization techniques, such as $k$-anonymity, $\ell$-diversity, and $t$-closeness, on ML fairness remain largely unexplored. In this work, we systematically audit the impact of anonymization techniques on ML fairness, evaluating both individual and group fairness. Our quantitative study reveals that anonymization can degrade group fairness metrics by up to four orders of magnitude. Conversely, similarity-based individual fairness metrics tend to improve under stronger anonymization, largely as a result of increased input homogeneity. By analyzing varying levels of anonymization across diverse privacy settings and data distributions, this study provides critical insights into the trade-offs between privacy, fairness, and utility, offering actionable guidelines for responsible AI development. Our code is publicly available at: https://github.com/hharcolezi/anonymity-impact-fairness.
@@ 1 Introduction
Ensuring fairness in ML involves mainly two key principles: (i) group fairness, which ensures that model predictions are consistent across demographic groups [13, 8], and (ii) individual fairness, which ensures that similar individuals receive similar treatment [18]. While anonymization techniques are effective in preserving privacy, they introduce transformations such as generalization or suppression, which may distort data distributions and induce unintended bias in ML models [3]. These alterations can inadvertently affect protected attributes, shift group distributions, and influence fairness metrics.
@@
Our contributions. In this paper, we present the first in-depth, systematic audit of the impact of three widely-used anonymization methods, namely, $k$-anonymity [38, 39], $\ell$-diversity [27], and $t$-closeness [26] on fairness in ML. While anonymization is widely used to meet privacy regulations such as GDPR, its downstream effects on fairness remain poorly understood. We fill this gap by examining their effects on group fairness metrics (e.g., equalized odds [23]), individual fairness metrics (e.g., similarity fairness [18]), while also examining the trade-offs with utility metrics (e.g., F1-score). Our findings highlight the nuanced interplay between anonymity, fairness, and utility, offering valuable insights and actionable guidance for practitioners working with anonymized datasets. In summary, the key contributions of this paper are:
@@ 2.1 Anonymization Methods
$k$-anonymity: A dataset $D$ satisfies $k$-anonymity with respect to a set of $QIs$, if and only if every equivalence class $E\subseteq D$ formed by the unique combinations of values in $QI$ contains at least $k$ records [38, 39]. Formally, $\forall E\subseteq D,\ |E|\geq k$, where $E$ is an equivalence class of records that share the same generalized values across $QI$. The $k$-anonymity property ensures that any individual represented in the dataset cannot be distinguished from at least $k-1$ other individuals based on the quasi-identifier attributes, thus reducing the risk of re-identification. Achieving $k$-anonymity typically involves generalization (replacing specific values with broader categories) or suppression of quasi-identifiers to create equivalence classes.
@@ 2.2 ML Fairness Metrics
Machine learning fairness [8] refers to the principle that ML models should produce predictions or decisions that are impartial and equitable across individuals or groups, particularly when protected attributes such as race, gender, or socioeconomic status are involved. This study employs a comprehensive evaluation of fairness using two primary categories of metrics: group fairness metrics (Section 2.2.1) and individual fairness metrics (Section 2.2.2).
@@
Equalized Odds Difference (EOD) [23] ensures that the model’s prediction is independent of the protected attribute, conditioned on the true label. It measures the disparity in true positive and false positive rates between two specific demographic groups defined by a protected attribute. Formally:
@@
Similarity Fairness (SF) assesses the degree to which a model treats similar inputs consistently. It evaluates the average variation in predictions for input pairs that are deemed similar based on a defined similarity metric. A lower SF value indicates better fairness, as it ensures consistent treatment of similar individuals. Formally:
$\text{SF}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}(x_{i})|}\sum_{x_{j}\in\mathcal{N}(x_{i})}\left|f(x_{i})-f(x_{j})\right|\cdot\text{sim}(x_{i},x_{j})\mathrm{,}$ (5)
where $\mathcal{N}(x_{i})$ represents the neighborhood of $x_{i}$, $\text{sim}(x_{i},x_{j})$ is the similarity score derived from the inverse of the distance, and $n$ is the total number of data points.
Neighborhood Consistency Fairness (NCF) evaluates the consistency of predictions within local neighborhoods, ensuring equitable treatment for instances with similar characteristics. A lower NCF value reflects better fairness, as it ensures that predictions are consistent for neighboring inputs. Formally:
$\text{NCF}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}(x_{i})|}\sum_{x_{j}\in\mathcal{N}(x_{i})}\mathbb{I}\left(f(x_{i})=f(x_{j})\right)\mathrm{,}$ (6)
where $\mathcal{N}(x_{i})$ is the neighborhood of $x_{i}$, $\mathbb{I}(\cdot)$ is the indicator function, which equals 1 if predictions for $x_{i}$ and $x_{j}$ are identical, and $n$ is the total number of data points.
@@ 3.1 Problem Formulation
While anonymization aims to protect individual privacy by transforming the data before learning, its downstream impact on fairness in ML is poorly understood. We therefore aim to systematically audit how different anonymization methods influence group and individual fairness outcomes, as well as predictive performance, across a range of ML models and datasets.
@@
  • $A\in\mathcal{A}$: Protected attribute(s) (e.g., race, gender), often binary ($A\in\{0,1\}$), with $A=1$ denoting the privileged group and $A=0$ the unprivileged group.
  • $X\in\mathcal{X}$: Non-sensitive, non-protected features (e.g., employment status, credit score).
  • $Y\in\{0,1\}$: Binary target variable (e.g., loan approval outcome), where $Y=1$ denotes a favorable decision.
Given this setting, a supervised ML model learns a predictive function $f:\mathcal{A}\times\mathcal{X}\rightarrow[0,1]$ such that $\hat{Y}=f(A,X)$ approximates the true label $Y$. In practice, the model is trained on a dataset $D^{\prime}$, which may differ from $D$ due to privacy constraints. Specifically, to preserve privacy, we consider an anonymization mechanism $\mathcal{A}$, which transforms the raw dataset $D$ into an anonymized dataset $D^{\prime}=\mathcal{A}(D)$. In this work, $\mathcal{A}$ represents one of the following methods: $k$-anonymity, $\ell$-diversity, or $t$-closeness, each of which enforces privacy by applying generalization and suppression over quasi-identifiers. We then train ML classifiers on both original ($D$) and anonymized ($D^{\prime}$) datasets, and compare them using performance (e.g., F1-score) and fairness metrics (Section 2.2).
@@ 3.2 Research Questions
  • RQ3) What is the impact of varying target distributions on the fairness of ML models? This research question examines how changes in the target distribution, specifically by varying the threshold used to binarize the target variable (see "Datasets" in Section 3.3), influence fairness metrics. Varying the target distribution affects the balance between positive and negative outcomes in the dataset, which can, in turn, affect fairness metrics between demographic groups. The experiments addressing this question are presented in Section 4.3.
  • RQ4) How does dataset size influence the relationship between anonymization and fairness in ML models? This question explores the role of dataset size in mediating the relationship between anonymization and fairness. By systematically varying the data fraction, we analyze how sample size influences the trade-offs between privacy, fairness, and utility. The experiments addressing this question are detailed in Section 4.4.
@@ 3.3 Experimental Setup
[…]
@@ 4 Results and Analysis
Following the methodology and experimental setup described in the previous Section 3, this section presents an analysis of the impact of anonymization methods on ML fairness. Due to space limitations, we present in the main paper the results obtained using the Adult dataset with gender as the protected attribute. Additional details and discussions, including results with race as the protected attribute and those for both Compas and ACSIncome datasets (evaluating both gender and race as protected attributes), are provided in Appendix B. However, it is important to emphasize that the discussions in this section are broadly applicable to the findings across both datasets and all protected attributes.
@@ 4.1 Impact of Anonymization on Fairness in ML
From Figure 1, it is evident that anonymization techniques negatively affect group fairness metrics such as SPD and EOD. For instance, as the parameter $k$ increases in $k$-anonymity, these fairness metrics tend to degrade, suggesting that stricter privacy constraints may amplify fairness disparities between demographic groups. Specifically, at $k=100$, the EOD metric increases from $\mathrm{EOD}=0.2$ (baseline $k=1$) to nearly $\mathrm{EOD}=0.8$ for both $k$-anonymity and $t$-closeness, representing an approximately fourfold increase. Similarly, the SPD metric rises from $\mathrm{SPD}=0.38$ (baseline $k=1$) to almost $\mathrm{SPD}=0.68$ (for $k$-anonymity and $t$-closeness). A similar trend is observed in Figure 6 for the Adult dataset with race as the protected attribute, where both SPD and EOD exhibit slight increases for some values of $k$. Furthermore, as shown in Figures 710, at least two group fairness metrics consistently degrade across the Compas and ACSIncome datasets, regardless of whether gender or race is used as the protected attribute. These findings highlight the broader implications of anonymization on fairness metrics across different datasets and protected attributes.
@@
Regarding individual fairness, Figure 1 reveals distinct trends across the three evaluated metrics: ALF, ASF, and NCF. ALF exhibits mixed behavior depending on the anonymization technique. For $k$-anonymity (and $0.55$-closeness), the metric tends to degrade for $2\leq k\leq 50$, indicating that stricter anonymity levels increase the sensitivity of model predictions to input changes. This suggests reduced stability in preserving Lipschitz fairness under $k$-anonymity. Conversely, for $\ell$-diversity and $0.45$-closeness, ALF shows improvement, demonstrating their potential to balance privacy and prediction stability for similar inputs. A similar trend is observed in Figure 6 for the Adult dataset with race as the protected attribute. In contrast, ASF generally worsens across all three anonymization methods in experiments with the ACSIncome dataset, irrespective of the protected attribute (i.e., see Figures 9 and 10).
These observations highlight the varied impact of anonymization on individual fairness metrics. While ALF demonstrates mixed trends, our results with all three datasets and protected attributes demonstrate that anonymization techniques positively affect both ASF and NCF individual fairness metrics (i.e., see Figures 1,  610). Specifically, ASF consistently improves with higher levels of anonymization across all three techniques. This improvement indicates that anonymization reduces variations in predictions for similar individuals, fostering more equitable treatment at the individual level. Notably, for $k\geq 10$, the ASF metric stabilizes at nearly zero, reflecting a substantial alignment in predictions for similar data points. Similarly, the NCF metric also benefits from anonymization, showing a clear decline in its values as $k$ increases. Lower NCF values imply fewer inconsistencies in predictions within local neighborhoods of similar inputs. For instance, $k$-anonymity demonstrates a significant drop in NCF at smaller $k$-values, with marginal improvements at higher levels of $k$. These trends can be attributed to the inherent differences in how ASF and NCF measure individual fairness. ASF captures the magnitude of prediction variations for similar inputs, penalizing even minor differences, whereas NCF focuses on exact consistency within local neighborhoods, relying on a binary indicator (see Eq. (6)). As anonymization increases, generalized attributes cause predictions to align more rapidly, leading to a faster decrease in ASF. In contrast, NCF, being less sensitive to small inconsistencies, decreases more gradually. Importantly, this smoothing effect and the resulting improvements in individual fairness metrics align with similar phenomena observed in differential privacy [18].
@@ 4.2 Detailed Analysis #1: Record Suppression Levels and Their Effect on Fairness in ML
[…]
From the results for both datasets and protected attributes presented in Figures 21115, the impact of suppression levels on fairness and utility metrics reveals nuanced patterns. For instance, the effect of suppression on group fairness metrics such as SPD, EOD, and MAD is more mixed. In some cases, increasing suppression levels slightly improve these metrics, likely due to the exclusion of outliers that disproportionately skew fairness. For example, SPD and EOD in the Adult dataset with race as the protected attribute (Figure 11) exhibit slight improvements at moderate suppression levels. However, in other cases, particularly for the ACSIncome dataset (Figures 14 and 15), the metrics remain stable or even worsen with higher suppression levels, suggesting that the removal of data points introduces disparities between demographic groups.

### v2

@@ Abstract
Machine learning (ML) algorithms are heavily based on the availability of training data, which, depending on the domain, often includes sensitive information about data providers. This raises critical privacy concerns. Anonymization techniques have emerged as a practical solution to address these issues by generalizing features or suppressing data to make it more difficult to accurately identify individuals. Although recent studies have shown that privacy-enhancing technologies can influence ML predictions across different subgroups, thus affecting fair decision-making, the specific effects of anonymization techniques, such as $k$-anonymity, $\ell$-diversity, and $t$-closeness, on ML fairness remain largely unexplored. In this work, we systematically audit the impact of anonymization techniques on ML fairness, evaluating both individual and group fairness. Our quantitative study reveals that anonymization can degrade group fairness metrics by up to fourfold. Conversely, similarity-based individual fairness metrics tend to improve under stronger anonymization, largely as a result of increased input homogeneity. By analyzing varying levels of anonymization across diverse privacy settings and data distributions, this study provides critical insights into the trade-offs between privacy, fairness, and utility, offering actionable guidelines for responsible AI development. Our code is publicly available at: https://github.com/hharcolezi/anonymity-impact-fairness.
[…]
3224
@@ 1 Introduction
Ensuring fairness in ML involves mainly two key principles: (i) group fairness, which ensures that model predictions are consistent across demographic groups [14, 9], and (ii) individual fairness, which ensures that similar individuals receive similar treatment [19]. While anonymization techniques are effective in preserving privacy, they introduce transformations such as generalization or suppression, which may distort data distributions and induce unintended bias in ML models [3]. These alterations can inadvertently affect protected attributes, shift group distributions, and influence fairness metrics. We hypothesize that such statistical alterations introduced to satisfy anonymization can significantly impact fairness metrics in ML.
@@
Our contributions. In this paper, we present the first in-depth, systematic audit of the impact of three widely-used anonymization methods, namely, $k$-anonymity [40, 41], $\ell$-diversity [29], and $t$-closeness [28] on fairness in ML. We analyze their effects on both group fairness metrics (e.g., equal opportunity [24]), individual fairness metrics (e.g., similarity fairness [19]), while also examining the trade-offs with utility metrics (e.g., F1-score). Our findings highlight the nuanced interplay between anonymity, fairness, and utility, offering valuable insights and actionable guidance for practitioners working with anonymized datasets.
In summary, the key contributions of this paper are:
@@ 2.1 Anonymization Methods
$k$-anonymity: A dataset $D$ satisfies $k$-anonymity with respect to a set of $QIs$, if and only if every equivalence class $E\subseteq D$ formed by the unique combinations of values in $QI$ contains at least $k$ records [40, 41]. Formally, $\forall E\subseteq D,\ |E|\geq k$, where $E$ is an equivalence class of records that share the same generalized values across $QI$. The $k$-anonymity property ensures that any individual represented in the dataset cannot be distinguished from at least k1k-1 other individuals based on the quasi-identifier attributes, thus reducing the risk of re-identification. Achieving $k$-anonymity typically involves generalization (replacing specific values with broader categories) or suppression of quasi-identifiers to create equivalence classes.
@@ 2.2 ML Fairness Metrics
Machine learning fairness [9] refers to the principle that ML models should produce predictions or decisions that are impartial and equitable across individuals or groups, particularly when protected attributes such as race, gender, or socioeconomic status are involved. This study employs a comprehensive evaluation of fairness using two primary categories of metrics: group fairness metrics (Section 2.2.1) and individual fairness metrics (Section 2.2.2). Their joint use allows us to examine the multifaceted impact of anonymization on both group-level disparities and individual-level consistency.
@@
Equal Opportunity Difference (EOD) [24] requires that predictions are independent of the protected attribute conditional on the positive class only (i.e., $\hat{Y}\perp A\mid Y{=}1$). It measures the disparity in true positive rates across groups. Formally:
@@
Similarity Fairness (SF) quantifies the local smoothness of the model: within each point’s neighborhood, it aggregates absolute prediction differences and weights them by the feature-space distance between pairs. This distance-weighted variant emphasizes consistency across the broader local region while de-emphasizing near-duplicate pairs. Lower SF values indicate smoother, more locally consistent predictions. Formally:
$\text{SF}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}(x_{i})|}\sum_{x_{j}\in\mathcal{N}(x_{i})}\left|f(x_{i})-f(x_{j})\right|\cdot\text{dist}(x_{i},x_{j})\mathrm{,}$ (5)
where $\mathcal{N}(x_{i})$ represents the neighborhood of $x_{i}$, $\text{dist}(x_{i},x_{j})$ is the distance between $x_{i}$ and $x_{j}$ in feature space, and $n$ is the total number of data points.
Neighborhood Consistency Fairness (NCF) evaluates the consistency of predictions within local neighborhoods, ensuring equitable treatment for instances with similar characteristics [43]. A lower NCF value reflects better fairness, as it ensures that predictions are consistent for neighboring inputs. Formally:
$\text{NCF}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{|\mathcal{N}(x_{i})|}\sum_{x_{j}\in\mathcal{N}(x_{i})}\mathbb{I}\left(f(x_{i})\neq f(x_{j})\right)\mathrm{,}$ (6)
where $\mathcal{N}(x_{i})$ is the neighborhood of $x_{i}$, $\mathbb{I}(\cdot)$ is the indicator function, which equals 1 if predictions for $x_{i}$ and $x_{j}$ are mismatched, and $n$ is the total number of data points.
@@ 3.1 Problem Formulation
We aim to systematically audit how different anonymization methods influence group and individual fairness outcomes, as well as predictive performance, across a range of ML models and datasets.
@@
  • $A\in\mathcal{A}$: Protected attribute(s) (e.g., race, gender), treated as quasi-identifiers for anonymization and explicitly included as input features. We adopt this approach in line with prior literature showing that excluding protected attributes (i.e., “fairness through unawareness”) does not prevent discriminatory outcomes [9, 26, 14]. We focus on binary protected attributes to ensure comparability across standard fairness metrics and to align with common preprocessing practices in the fairness and privacy literature [42].
  • $X\in\mathcal{X}$: Non-sensitive, non-protected features (e.g., employment status, credit score), which are also treated as quasi-identifiers, thus, subject to generalization or suppression.
  • $Y\in\{0,1\}$: Binary target variable (e.g., loan approval outcome), where $Y=1$ denotes a favorable decision. We focus on binary targets for interpretability and comparability across fairness metrics, consistent with standard practices in the fair-ML [43, 9] and privacy-fairness literature [8, 5, 13, 16, 31].
Given this setup, a supervised ML model learns a predictive function $f:\mathcal{A}\times\mathcal{X}\rightarrow[0,1]$ such that $\hat{Y}=f(A,X)$ approximates the true label $Y$. In practice, the model is trained on a dataset $D^{\prime}$, which may differ from $D$ due to privacy constraints. Specifically, to preserve privacy, we consider an anonymization mechanism $\mathcal{M}$, which transforms the raw dataset $D$ into an anonymized dataset $D^{\prime}=\mathcal{M}(D)$. In this work, $\mathcal{M}$ represents one of the following methods: $k$-anonymity, $\ell$-diversity, or $t$-closeness, each of which enforces privacy by applying generalization and suppression over quasi-identifiers. We then train ML classifiers on both original ($D$) and anonymized ($D^{\prime}$) datasets, and compare them using performance (e.g., F1-score) and fairness metrics (Section 2.2).
@@ 3.2 Research Questions
  • RQ3) What is the impact of varying target distributions on the fairness of ML models? This research question examines how changes in the target distribution, specifically by varying the threshold used to binarize the target variable (see "Datasets" in Section 3.3), influence fairness metrics. In many real-world settings, continuous outcomes (e.g., income, risk score) are binarized to support binary decision-making processes such as loan approvals, benefit eligibility, or pretrial release decisions. However, the choice of threshold can significantly affect the proportion of positive versus negative labels, altering group-specific outcome rates and thereby impacting fairness metrics. The experiments addressing this question are presented in Section 4.3.
  • RQ4) How does dataset size influence the relationship between anonymization and fairness in ML models? This question explores the role of dataset size in mediating the relationship between anonymization and fairness. In practice, smaller datasets may be more vulnerable to the distortions introduced by anonymization, such as over-generalization or excessive suppression, which could disproportionately affect minority groups or rare feature combinations. By systematically varying the data fraction, we analyze how sample size influences the trade-offs between privacy, fairness, and utility. The experiments addressing this question are detailed in Section 4.4.
@@ 3.3 Experimental Setup
Anonymization parameters: The anonymization parameters are varied as follows: $k\in\{1,2,\ldots,9,10,25,50,75,100\}$ for $k$-anonymity, $\ell=2$ for $\ell$-diversity (binary target), and $t\in\{0.45,0.50,0.55\}$ for $t$-closeness [^1] The values of $t$ specify upper bounds on the Earth Mover’s Distance between the distribution of the sensitive attribute in the entire dataset and its distribution within each equivalence class.. Note that for $k=1$, no anonymity is satisfied, which serves as the non-private baseline. Unless otherwise mentioned, the allowed record suppression level is fixed at $\textsf{supp\_level}=20\%$ [^2] Our attempts with lower suppression levels (0%-15%) did not achieve anonymization for high $k$ values because we operated under a worst-case assumption where all attributes except the target were treated as QIDs.. The anonymization methods were implemented with Anjana [39]. The same generalization levels applied to the training set are replicated for the test set to ensure consistency and prevent discrepancies between training and evaluation.
@@ 4 Results and Analysis
Following the methodology and experimental setup described in the previous Section 3, this section presents an analysis of the impact of anonymization methods on ML fairness. Due to space limitations, we present in the main paper the results obtained using the Adult dataset with gender as the protected attribute. Additional details and discussions, including results with race as the protected attribute and those for both Compas and ACSIncome datasets (evaluating both gender and race as protected attributes), are provided in Appendix B. However, it is important to emphasize that the discussions in this section are broadly applicable to the findings across all datasets and all protected attributes.
@@ 4.1 Impact of Anonymization on Fairness in ML
From Figure 1, it is evident that anonymization techniques negatively affect group fairness metrics such as SPD and EOD. For instance, as the parameter $k$ increases in $k$-anonymity, these fairness metrics tend to degrade, suggesting that stricter privacy constraints may amplify fairness disparities between demographic groups. Specifically, at $k=100$, the EOD metric increases from $\mathrm{EOD}=0.2$ (baseline $k=1$) to nearly $\mathrm{EOD}=0.8$ for both $k$-anonymity and $t$-closeness, representing an approximately fourfold increase. Similarly, the SPD metric rises from $\mathrm{SPD}=0.38$ (baseline $k=1$) to almost $\mathrm{SPD}=0.68$ (for $k$-anonymity and $t$-closeness). However, this degradation is not always monotonic: we observe fluctuations in fairness metrics at intermediate privacy levels, revealing a more complex, non-linear relationship between anonymity strength and fairness outcomes. A similar trend is observed in Figure 6 for the Adult dataset with race as the protected attribute, where both SPD and EOD exhibit slight increases for some values of $k$, but not in a strictly increasing manner. Furthermore, as shown in Figures 710, at least two group fairness metrics degrade across the Compas and ACSIncome datasets, regardless of whether gender or race is used as the protected attribute. These findings highlight the broader implications of anonymization on fairness metrics across different datasets and protected attributes.
Individual fairness.
@@
Figure 1 also reveals diverse trends across the three evaluated metrics. ALF shows non-monotonic and method-dependent behavior. Under $k$-anonymity (and $0.55$-closeness), ALF generally worsens for $2\leq k\leq 50$, indicating that stricter anonymity levels increase model sensitivity to small input changes, suggesting reduced robustness in preserving Lipschitz fairness. In contrast, under $\ell$-diversity and $0.45$-closeness, ALF improves, highlighting their potential to better preserve prediction stability among similar individuals. A similar pattern is observed in Figure 6 for the Adult dataset with race as the protected attribute.
In contrast, both ASF and NCF consistently improve across all three anonymization methods, datasets, and protected attributes (i.e., see Figure 1 and Figures 610). Specifically, ASF shows substantial gains with increasing levels of anonymization, reflecting reduced prediction variability among similar individuals and promoting more equitable treatment at the individual level. For $k\geq 10$, ASF values stabilize near zero, indicating high alignment in predictions across local neighborhoods. NCF also improves with anonymization, though more gradually, as it relies on exact prediction matches. For instance, $k$-anonymity achieves a sharp drop in NCF at lower $k$-values, with smaller gains at higher levels. These trends reflect the inherent differences between ASF and NCF: while ASF penalizes even small deviations, NCF uses a binary notion of consistency (see Eq. (6)). Overall, the smoothing effect introduced by generalization appears to benefit individual fairness, consistent with similar effects observed under differential privacy [19].
Utility.
@@ 4.2 Detailed Analysis #1: Record Suppression Levels and Their Effect on Fairness in ML
Following the findings of Section 4.1, to answer RQ2 from Section 3.2, we now investigate the impact of the allowed record suppression level (supp_level) on fairness and utility in ML. Record suppression, which involves removing rows during anonymization, can exclude outliers from the dataset, potentially worsening fairness metrics as certain demographic groups may be disproportionately impacted. To understand this effect, we vary the suppression level ($\texttt{supp\_level}\in\{10,20,30,40,50\}$) and evaluate its impact on both group and individual fairness metrics, as well as utility. For the remainder of this analysis, the anonymity parameters are fixed at $k=10$ for $k$-anonymity and $t=0.5$ for $t$-closeness. The choice of $k=10$ represents a moderate level of anonymity, balancing privacy protection and data utility. Similarly, $t=0.5$ ensures a reasonable level of distributional closeness under $t$-closeness, reflecting practical settings often used in real-world applications. Fixing these parameters allows for a focused evaluation of how varying suppression levels impact fairness and utility, independent of additional variability introduced by the anonymity parameters.
Figure 2: Effect of allowed record suppression level ($\texttt{supp\_level}\in\{10,20,30,40,50\}$) in anonymization techniques ($10$-anonymity, 22-diversity, $0.5$-closeness) on group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. The results are derived from the Adult dataset, using gender as the protected attribute for fairness evaluation.
From the results for all datasets and protected attributes presented in Figure 2 and Figures 1115, the impact of suppression levels on fairness and utility metrics reveals nuanced patterns. For instance, the effect of suppression on group fairness metrics such as SPD, EOD, and MAD is more mixed. In some cases, increasing suppression levels slightly improve these metrics, likely due to the exclusion of outliers that disproportionately skew fairness. For example, SPD and EOD in the Adult dataset with race as the protected attribute (Figure 11) exhibit slight improvements at moderate suppression levels. However, in other cases, particularly for the ACSIncome dataset (Figures 14 and 15), the metrics remain stable or even worsen with higher suppression levels, suggesting that the removal of data points introduces disparities between demographic groups.